Recent Log Entries

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Log Number: 3876711Logbook(s): Body: 

Day #157, Week 32 Day 2, Wrap-Up:

AIPINJ --
* Photocathode puck #9 SLSP-5756-4 (CREX) heated in the morning, activated in the afternoon. 3876{601, 707}
* Yellow laser box cable entry re-sealed with black tape. 3876616
* New, EES-IC standard, collision circuit installed for 0I04 (viewer, master slit, chopping apertures, central plug) and verified good. During install, found MS out limit-switch bad; replaced. 0I04 viewer MAC valve found bad; replaced, along with readjusting viewer's out limit-switch.

COMMENT ON Install C-100 style vacuum controls and cabling in zone NL05

Log Number: 3876710Logbook(s): Body: 

Waveguide ion pumps (VIP1L051-VIP1L058) connected to
service cables. Pumps powered from rack supplies, with live
EPICS readbacks. Pressure readbacks for all waveguide IPs
in low to mid e^-10s.

Frank Humphry has the beamline pump "1L05B" on a local
supply. (Ref. Task 22374).

Follow-up Re: UITF Loss of SF6 ALARM

Log Number: 3876709Logbook(s): Body: 

The Glassman gun HVPS interlock chain chassis has an SF6 pressure interlock which is connected to an electronic sensor pressure gauge in the HVPS SF6 tank in the UITF vault. This electronic pressure sensor has a low limit set point, which was set to 50 psig. Because the HVPS tank has a very small leak, the pressure has been dropping slowly over months and today it reached 50.1 psig which triggered the buzzing alarm.

Photocathode SLSP-5756-4 activated

Log Number: 3876707Logbook(s): Entry Makers: gramesBody: 

After cooling Puck #9 SLSP-5756-4 to 29 C during the day shift activated with Cs (4.5A) and NF3 (+5-10 nA) using 773nm diode, 5 mm mask and bias=5. Turn on came a little early ~15 min, completed 15 yo-yo's w/ adjustments to add ND filter and finished on ~1 uA FS.

Stripchart was mis-behaving, so a very touch-feely activation through first 5 yo-yo's, need to address that, maybe picoammeter, bias seemed okay but worth a look.

UITF Loss of SF6 ALARM

Log Number: 3876704Logbook(s): Body: 

The UITF
"LOSS OF SF6 ALARM" panel is lit and
the very loud alarm can be heard throughout the Test Lab high bay.

Panel is in rack # ITF 21, Gun HV and Laser on the UITF mezzanine.
I did not see a reset button or list of immediate actions.

COMMENT ON DC Power Support for Injector Upgrade

Log Number: 3876699Logbook(s): Body: 

All the magnets cable for IN02B11 trim rack has been verified
by ohm in out the cables at the junction box and at the magnet
connection leads. All of the IN02B11 magnets are hookup at
the magnets terminal strips.

We will continue with IN02B10 trim rack tomorrow.

Diggs, Kissinger, Richardson

COMMENT ON *Vacuum Activities CEBAF 03/03/20 through

Log Number: 3876695Logbook(s): Body: 

Girder's 1L26 and 1L27 + Dummy are 2A alignment and
readied for Hook-up.

Arc 7 Beamline is being Pumped down will get a leak check
(hopefully) tomorrow.

Appendage VIP added to VIP4S00A will Pumpdown and Leak
Check tomorrow.

Injector Current Monitoring Screen Updated

Log Number: 3876693Logbook(s): Entry Makers: moserBody: 

Added A1-A4 aperture hole size PVs. Added MS insertion status readback. Added link to A1 remote controls screen.

These updates are in addition to the ones described in the linked log.

MBH1I07H setpoint change

Log Number: 3876691Logbook(s): Entry Makers: hoflerBody: 

I was moving a corrector screen on my desktop, and instead of moving it , I somehow grabbed the slider for MBH1I07H.BDL and changed the setpoint for MBH1I07H. I put it back to 0.

COMMENT ON C75-01 Cryomodule installation into 1L05

Log Number: 3876687Logbook(s): Body: 

C75-01 cryomodule was installed in 1L05 at 12:30. The transport wheels were removed. The module alignment will start in the morning.
Leo was notified about hooking up the beamline and waveguide ion pumps.

Phone Outage Saturday

Log Number: 3876680Logbook(s): Body: 

Campus-wide Phone Outage Scheduled

INTERNAL MEMO

May 11, 2021
Sent on behalf of the CST Division

SUBJECT: Campus-wide Phone Outage Scheduled
CONTACT: CST Division Help Desk, helpdesk@jlab.org, x7155

Pages